# Stimulating Imagination: Towards General-purpose Object Rearrangement

### v1

@@ 3.3 GPT-assisted object rearrangement data generation
The available amount of public 3D object rearrangement data is limited, especially the data containing low-level rearrangement instructions. In this work, we develop an automatic pipeline for generating high-quality rearrangement-instruction pairs. Each instance comprises an initial scene, a goal scene after the rearrangement and a corresponding instruction. A total of 40,000 stable and collision-free instances are generated in the PyBullet physics simulator [9], rendered by OpenGL [1]. The simulated objects are randomly selected from the popular ShapeNetSem [6] (specifically ShapeNetSem [33]) dataset. We collect various 581 objects from 30 categories to ensure diversity.
@@
The entire generation pipeline consists of three steps: (1) pre-processing metadata to obtain well-constructed and realistic simulated objects; (2) randomly selecting the reference, moving and irrelevant objects, namely $O_{r}$, $O_{m}$ and $\textbf{O}\setminus\{O_{r},O_{m}\}$, loading them to PyBullet to obtain the initial and goal scenes, then filtering out physically-unrealistic ones; (3) using GPT-4 to generate rearrangement language instructions corresponding to the transition from the initial scene to the goal scene, based on the object and scene information. The difficulty lies in the time-consuming and cumbersome data collection process, as well as the limited capability of precise (fine-grained) spatial understanding and reasoning in existing models, even GPT-4.
@@
Pre-processing. The metadata in ShapeNetSem needs to be pre-processed by scaling and translation, because the object models may have unrealistic sizes and the centroids of objects may not be aligned with the origin of the point-cloud coordinate system. The scaling factors are obtained with GPT-4, e.g., asking GPT-4 about the typical size of a cellphone and accordingly scaling the object model.
@@
Scene generation. We categorize the data generation into multiple scenarios according to the relative spatial relationship between the moving and reference objects, such as left, right, front, behind, on, between, etc. We randomly select the object set O and the scenario to generate scenes. For the initial scene, we place all the objects with random positions in PyBullet, wait for them to settle into stability and record their poses. For the goal scene, we replace the reference and irrelevant objects with the recorded poses and then load $O_{m}$. Its pose is randomly sampled within a region determined by the $O_{r}$ and the scenario. For example, in the coordinate system with $O_{r}$ as the origin, “left” refers to the region $\{(x,y)\in\mathbb{R}^{2}\,|\,x/\sqrt{x^{2}+y^{2}}<-\delta,|y|/\sqrt{x^{2}+y^{2}}<\delta\}$, where $\delta$ is a hyper-parameter. The physical validity is verified in two aspects: the stability is measured by the angular and linear velocities in the physics engine, the collision is determined by checking whether the positions of $\textbf{O}\setminus\{O_{m}\}$ has any displacement.
@@
Instruction generation. Inspired by previous works [22, 21], we use GPT4 to generate the instruction in natural language form, given spatial coordinates and object information (e.g., RGB value and size) of scenes. We observe that it is essential to provide detailed descriptions of spatial and object information, otherwise GPT may not be able to understand the spatial transition or determine whether the placement is reasonable.
@@
An interesting observation is that we have tried an end-to-end instruction generation approach by directly prompting GPT4 the rendered RGB images of the initial and final scenes. But the generated language instructions are of low accuracy. We attribute this to two main reasons: (1) the absence of lifelike qualities in the simulated objects impedes GPT4’s ability to recognize them accurately, (2) deducing fine-grained 3D spatial relationships from RGB images, which usually requires considering occlusion and perspective effects, is challenging for existing LLMs [43] (even the powerful GPT4). We will continue to explore this topic in future work.
4 Experiments
@@
The goal of the experiments is to evaluate the efficacy of SPORT in the object rearrangement task, especially in the ability of generalization, 3D spatial reasoning and precise instruction following. To this end, we need to answer the following questions:
@@
The experiments are then conducted both in simulation and real-world environments. The experimental details and results are reported in the following two sections.
Table 1: (Performance in simulation) “Pose Accuracy” refers to the accuracy of whether the objects are correctly recognized and whether the generated poses of objects satisfy the instruction. Six scenarios and summary data are reported. “Physical Realism” refers to the accuracy of whether the rearrangement is physically-realistic. The “Overall Success” is achieved only when both requirements are met. “SPORT*” indicates the model using object masks from the simulation environment.
Pose Accuracy Physical Overall
On Between Front Behind Left Right ALL Realism Success
SPORT 51.18 35.75 63.95 63.69 63.27 67.19 59.64 70.48 46.19
SPORT* 76.63 91.70 91.22 90.20 90.00 89.95 87.80 76.40 69.49
4.1 Simulation experiments
Setup To fully validate the generalization ability of SPORT, we conduct a cross-dataset evaluation. Unlike the objects in training data (from ShapeNetSem), the simulated objects for testing are collected from Google Scanned Objects [11]. A total of 77 object models from 37 novel categories are randomly selected. We use PyBullet [9] as the physics simulator and OpenGL [1] as the appearance render. For each testing sample, the involved objects are randomly sampled. The scene and the rearrangement instruction are generated following the pipeline in subsection 3.3.
We use the success rate as the evaluation metric as in previous works [23]. There are three aspects to consider: given a command, the model should be able to recognize the involved objects, place them to correct positions, and the rearrangement is physically-realistic. We systematically measure whether the placed positions of objects satisfy the command, with similar rules for assessing spatial relationships in subsection 3.3. For example, given the command “put $O_{m}$ to the left of $O_{r}$”, in the coordinate system with $O_{r}$ as the origin, the coordinates $(x,y)$ of $O_{m}$ should satisfy $x/\sqrt{x^{2}+y^{2}}<-\delta^{\prime},|y|/\sqrt{x^{2}+y^{2}}<\delta^{\prime}$. As for the assessment of physical validity, we continuously place objects in simulation based on the estimated poses, checking the collision and stability. A rearrangement is considered as correct only when all these aspects are satisfactory.
@@ 4.1 Simulation experiments
The results are listed in Table 1. Whether a testing sample is classified as correct depends on three aspects: the objects are correctly recognized, the generated poses of objects satisfy the instruction, and the rearrangement is physically-realistic. “Pose accuracy” refers to the accuracy of considering the first two aspects, “Physical Realism” focuses on the last one, and “Overall Success” is the overall success rate considering all three aspects. Six scenarios (corresponding to six spatial relationships) are conducted for evaluation, namely left, right, front, behind, on and between.
As shown in Table 1, SPORT achieves an overall success rate of $46.19\%$ on the simulation testing set. The result is acceptable due to the challenging experimental setting: the testing and training data are collected from different datasets (the objects are totally different), and the final states of objects should stable and collision-free. However, we want to explore more, especially given the observation that Physical Realism achieve higher accuracy than Pose Accuracy, which is quite unusual.
We notice that LISA fails a lot on images rendered in simulation. On simulation-rendered images, the completeness at the edges of the object segmentation masks produced by LISA are not sufficient. As a result, the quality of the resultant 3D point clouds is often poor. We attribute this to the substantial domain differences between the simulation images and the real-world images used in LISA’s training set, resulting in limited model generalization. However, LISA can indeed localize target objects even requiring complex reasoning in real-world scenarios. We have conducted related experiments, please refer to subsection 3 for the details.
According to the above observation, we want to know the performance of SPORT if the required objects can be successfully obtained (since LISA can achieve this on real-world images). The results are listed in the second row, by using object masks directly from simulation data. One can see that SPORT achieves convincing performance, $87.8\%$ on Pose Accuracy and $69.49\%$ on Overall Success.
Finally, one can see that our approach demonstrates a certain degree of effectiveness in generating physically-realistic poses. The success rate of Physical Realism achieves $76.40\%$, assessed by using the Pybullet simulator. Despite the progress, this particular ability indeed needs further enhancement. Its accuracy exhibits a disparity relative to Pose Accuracy. We leave the exploration as further work.
Table 2: (Ablation study) The effects of two training configurations are discussed. The base model is SPORT*. Please refer to Table 1 for details. Six scenarios are included. The overall success rates (including physical validation) for each one and ALL are reported.
On Between Front Behind Left Right ALL
SPORT* 44.84 62.66 78.82 78.63 77.65 77.51 69.49
Poses trainable 46.80 44.58 75.39 79.60 78.57 74.25 65.59 (-3.90)
BERT trainable 16.53 57.50 20.47 20.74 21.76 20.15 27.65 (-41.84)
Table 3: (Ablation study: the effect of the scale of training data) The base model is SPORT*. The overall success rates (including physical validation) for each scenario and ALL are reported.
On Between Front Behind Left Right ALL
100% data 44.84 62.66 78.82 78.63 77.65 77.51 69.49
50% data 40.16 52.72 74.09 70.75 72.55 68.95 62.42 (-7.07)
25% data 36.51 54.13 66.22 64.06 61.33 67.40 58.12 (-11.37)
10% data 17.66 42.68 57.37 57.43 56.01 60.63 46.73 (-22.76)
@@
In this part, we perform ablation studies to synthetically analyze the proposed SPORT.
Estimating Poses of Reference and Irrelevant Objects. In our approach, the poses of the reference and irrelevant objects are fixed when training the diffusion model. We conduct an experiment to explore the impact of such a design, i.e., comparing the performances of fixing poses versus not fixing them. As shown in Table 2, fixing the poses significantly aids the model in learning relative positional relationships, effectively improving the success rate, from $65.59\%$ to $69.49\%$. It helps the model identify the reference points and possible collisions from the beginning of training, focusing on the learning of replacing the target object according to the command.
Making the instruction encoder trainable. We train the instruction encoder in the 3D goal estimator to study its effect on the performance. The performance comparison of training versus not training the encoder is listed in Table 2. One can observe that freezing the text encoder achieves better results. This is really an interest observation, since normally end-to-end tuning is a standard procedure. We speculate that the reason is that pre-trained BERT is already good enough for text understanding, while the amount of our collected data is insufficient for training all the modules in diffusion model. Similar observations can be found in [45].
Training on different scales of data. To investigate the impact of different scales of training data on model performance, we design a series of comparative experiments. In these experiments, the model is trained with 10%, 25%, and 50% of the entire dataset, while the network architectures and other training configurations stay the same. The results, as presented in Table 3, indicate a clear trend: as the volume of training data increases, the model’s success rate correspondingly improves, but the rate of improvement diminishes with larger data volumes. The diminishing return suggests that further increasing the training data volume yields only marginal benefits. That is, further expansion of the training dataset is unlikely to provide significant additional benefits for our task. This observation supports the conclusion that the goal estimator does not require massive data for training.
4.2 Real-world experiments
Figure 3: Examples of object rearrangement in the real world. The initial scenes are captured with a RGB-D camera, and then the model predicts the goal scene following the instruction.
Setup We collect several real-world scenes, captured by an Intel RealSense D435i RGB-D camera. Each scene includes several common objects placed on a table. The objects include various fruits, food, potato chips, tableware, etc. Challenging evaluations can be conducted that require complex spatial reasoning involving multiple yet similar objects.
@@ 4.2 Real-world experiments
SuSIE [2] leverages a diffusion model to “edit” the image of current scene to generate the intermediate subgoal image based on instructions. A low-level policy then executes the actions to reach the subgoal. Object rearrangement can be done by alternating this loop. The key is to implement the diffusion model with InstructPix2Pix [4], a powerful image-editing model pre-trained on Internet-scale data.
AVDC [19] uses an image diffusion model to synthesize a video of imagined execution of the rearrangement process. The underlying assumption is that the video generation model is a “world model” being capable of predicting the future. We utilize the public AVDC model trained on Bridge dataset [12] as the competitor, which is a real-world video dataset.
@@
Figure 3 illustrates the comparisons. One can observe that SPORT performs significantly better than the competitors. It can “imagine” the object positions strictly following the command, performing the replacement directly in 3D space. AVDC produces fuzzy and less-realistic videos, and often fails to generate correct goal images. Actually we cannot fully reproduce the results of SuSIE, since we do not have the authors’ robot setup. Nevertheless, we follow the code of SuSIE to produce goal images by repeating the image generation several times. Some issues can be observed to some extent, e.g., hallucination and loss of details. Moreover, both of these methods would hallucinate a robot or a human arm, originated from training data, which we believe may affect generalization.
5 Conclusions
In this work, we demonstrate the potential for achieving general-purpose object rearrangement by combining a pre-trained large vision model with a diffusion-based 3D pose estimation model. Given an instruction in natural language format, an LLM is used to identify the objects to be moved and to be reference. Given RGB-D images, we utilize an LLM-enhanced image segmentation model to segment required objects and then obtain their 3D point clouds. Based on these results, a diffusion-based 3D pose estimation model can follow precise low-level instructions to achieve physically-realistic position predictions. By establishing a GPT-assisted pipeline from a 3D perspective, a high-quality dataset for the object rearrangement task is generated. The results from both simulation and real-world experiments demonstrate the effectiveness of our approach. The model trained with simulation data can seamlessly transfer to real-world scenarios, achieving promising performance.

### v2

@@ III-C GPT-assisted Data Generation
The public available data for “place something relative to something” is limited. In this work, an automatic pipeline is proposed for collecting high-quality placement-instruction pairs. Each instance comprises an initial scene, a goal scene after the rearrangement and a corresponding instruction. A total of 40K stable and collision-free instances are generated in the PyBullet physics simulator [31], rendered by OpenGL [32]. The simulated objects are randomly selected from the popular ShapeNetSem [33, 34] dataset. We collect various 581 objects from 30 categories to ensure diversity.
@@
The pipeline consists of three steps: (1) pre-processing metadata to obtain well-constructed and realistic simulated objects; (2) randomly selecting the reference, movable and irrelevant objects, namely $O_{r}$, $O_{m}$ and $\textbf{O}\setminus\{O_{r},O_{m}\}$, loading them in PyBullet to obtain the initial and goal scenes (also filtering out physically-unrealistic ones); (3) using GPT-4 to generate language instructions corresponding to the transition from the initial to the goal scenes. The difficulty lies in the time-consuming and cumbersome data collection process, as well as the limited capability of precise (fine-grained) spatial understanding in existing models, even GPT-4.
@@
Pre-processing. The object models in ShapeNetSem may have unrealistic sizes and the centroids of them may not be aligned with the origin of the point-cloud coordinate system. Thus scaling and translation are required. The scaling can be done with GPT-4, e.g., asking GPT-4 about the typical size of a cellphone and accordingly scaling the object model.
@@
Scene generation. We build multiple scenarios according to the relative spatial relationship between the movable and reference objects, such as left, right, between, etc. The object set O and the scenario are randomly selected to generate scenes. For the initial scene, we place all objects with random positions in PyBullet, recording the poses once they are stabilized. For the goal scene, we replace the reference and irrelevant objects with the recorded poses and then load $O_{m}$. Its pose is randomly sampled within a region. For example, in the coordinate system with $O_{r}$ as the origin, “left” refers to the region $\{(x,y)\in\mathbb{R}^{2}\,|\,x/\sqrt{x^{2}+y^{2}}<-\delta,|y|/\sqrt{x^{2}+y^{2}}<\delta\}$ ($\delta$ is a hyper-parameter). The physical validity refers to stability and collision. They are measured by angular and linear velocities in physics engine, and checking whether the positions of $\textbf{O}\setminus\{O_{m}\}$ has any displacement, respectively.
@@
Instruction generation. Inspired by previous works [25], we use GPT4 to generate natural language instructions, given the spatial coordinates and object information (e.g., RGB value and size) of scenes. We observe that without such rich information, GPT may probably fail to understand the spatial transition or determine whether the placement is reasonable.
@@
Another interesting observation is that we have tried an end-to-end instruction generation approach: directly prompting GPT4 the rendered RGB images of the initial and final scenes. However, the generated language instructions are of low accuracy. We attribute this to two main reasons: (1) the absence of lifelike qualities in simulated objects impedes GPT4’s recognition ability; (2) deducing fine-grained 3D spatial relationships from only RGB images is challenging for existing LLMs [35] (even the powerful GPT4).
IV Experiments
@@
The goal of the experiments is to evaluate the efficacy of SPORT on the task of place something relative to something, especially in the ability of generalization, 3D spatial reasoning and precise instruction following. To this end, we need to answer the following questions:
@@
The evaluation are performed in both simulation and real-world environments.
IV-A Simulation Experiments
Setup A cross-dataset evaluation is conducted to fully validate the generalization ability of SPORT. Different from the training data (ShapeNetSem), the simulated objects for testing are collected from Google Scanned Objects [36]. A total of 77 object models from 37 novel categories are randomly selected. We use PyBullet [31] as the physics simulator and OpenGL [32] as the appearance render. 4K testing instances (scene and instruction) are generated following the pipeline in subsection III-C, within which the involved objects are randomly sampled.
As in previous works [6], evaluation metric is the success rate. There are three aspects to consider: given a command, the model should recognize the involved objects, place them to correct positions, and the rearrangement is physically realistic. We systematically measure whether the placed positions of objects satisfy the command, with similar rules for assessing spatial relationships in subsection III-C. To assess physical validity, we continuously place objects in simulation based on the estimated poses, checking the collision and stability. A placement is considered as correct ONLY when all these aspects are satisfactory.
@@ IV-A Simulation Experiments
TABLE I: (Performance in simulation) All testing objects are unseen to the model. “Pose Accuracy” refers to the accuracy of whether the final object poses satisfy the instruction. “Physical Realism” refers to the accuracy of whether the placement is physically realistic. “Overall Success” is achieved only when the above two requirements are met.
Pose Accuracy Physical Realism Overall Success
SPORT 59.64 70.48 46.19
SPORT* 87.80 76.40 69.49
Poses-train 83.46 77.68 65.59
BERT-train 36.38 75.04 27.65
50% data 80.48 72.16 62.42
25% data 76.17 71.59 58.12
10% data 64.89 63.17 46.73
@@
Six scenarios (corresponding to six spatial relationships) are conducted for evaluation. The results are listed in Table I. Whether a testing sample is classified as correct depends on three aspects: the objects are recognized, the generated object poses satisfy the instruction, and the rearrangement is physically-realistic. “Pose Accuracy” refers to the accuracy of the first two aspects, “Physical Realism” focuses on the last one, and “Overall Success” is the overall success rate considering all three aspects. Moreover, the “Pose Accuracy” for each scenario are shown in Fig. 3.
It can be seen that SPORT achieves an overall success rate of $46.19\%$. The result is acceptable due to the challenging experimental setting: the testing and training data are from different datasets (the 77 testing objects are unseen to the model), and the final states of objects should be physically valid. However, we want to explore more, especially given the observation that Physical Realism achieve higher accuracy than Pose Accuracy, which is quite unusual.
We notice that LISA fails a lot on simulation-rendered images. The main reason is the substantial domain differences between the simulation images and the real-world images used in LISA’s training set. On rendered images, the completeness of the object segmentation masks produced by LISA is lacking, and thus the quality of the resultant 3D point clouds is often poor. However, LISA can indeed localize target objects even requiring complex reasoning in real-world scenarios. We have conducted real-world experiments, please refer to subsection 4 for the details.
According to this observation, we want to know the performance if the required objects can be successfully obtained in simulation (because LISA can achieve this on real-world images), by using object masks directly from simulation data. The results are listed in the second row in Table I, abbreviated as SPORT*. It achieves convincing performance, $87.8\%$ on Pose Accuracy and $69.49\%$ on Overall Success.
Figure 3: Performance of six specific scenarios in simulation.
Finally, our method demonstrates a certain degree of effectiveness in generating physically-realistic poses. This is a challenging task, requiring the generative model to directly estimate stable, collision-free and physically feasible poses. SPORT* achieves $76.40\%$ on Physical Realism, assessed by using the Pybullet simulator.
@@
The baseline of ablation studies is SPORT*.
Estimating Poses of Reference and Irrelevant Objects. In our approach, the poses of the reference and irrelevant objects are fixed when training the diffusion model. An exper-iment is conducted to explore the impact of this design. The performance of not fixing these poses is listed in the third row in Table I. The success rate decreases from $69.49\%$ to $65.59\%$. Such a design helps the model identify the reference points and possible collisions, focusing on the learning of replacing the target object according to the command.
Making the instruction encoder trainable. We train BERT in the goal pose estimator to study its effect on the performance. As listed in the fourth row in Table I, the performance decreases significantly. This is really interesting since end-to-end tuning is quite standard. The probable reason is that our collected data is insufficient to train all network modules, while pre-trained BERT is already good enough for text understanding. Similar observations can be found in [37].
Training on different scales of data. The effect of different scales of training data on model performance is studied here. Three models, trained with 10%, 25%, and 50% of the entire dataset, are compared. The results in Table I indicate a clear trend: as the volume of training data increases, the model’s performance correspondingly improves. We do not further study the scaling law due to the extensive resources and time required. We leave it for future work.
IV-B Real-world Experiments
Figure 4: Results for the five tasks on the real-world planar scene. The first column shows initial scenes captured with a RGB-D camera. Our model is trained in simulation and tested directly on these real-world cases.
Setup We collect 10 real-world scenes, captured by an Intel RealSense D435i RGB-D camera. Accomplishing these real-world tasks requires complex spatial reasoning involving multiple, yet similar, objects. Moreover, the last five scenes include objects placed on a 3D shelf, which is very different from the training scenes and thus more challenging.
@@ IV-B Real-world Experiments
SuSIE [2] and AVDC [38] both leverage diffusion models (e.g., [39]) to synthesize the imagined execution process of placement. The former generates subgoal images in an “edit” manner while the latter synthesizes a video directly.
Figure 5: Results for the five tasks on the real-world 3D shelf scene. This kind of task that requires placing objects within 3D structures is more challenging because the model has not encountered it during training.
@@
Fig. 4 and Fig. 5 illustrate the results. SPORT performs significantly better than the competitors. It can “imagine” the object positions strictly following the command, performing the rearrangement directly in 3D space. For example, SPORT can identify Budweiser beer from canned drinks and predict its accurate pose, even if the placement needs to be performed within a 3D structure (the first row in Fig. 5).
In comparison, AVDC produces fuzzy and less-realistic videos, and often fails to generate correct goal images. As for SuSIE, actually the results cannnot be fully reproduced since we do not have the authors’ robot setup. Nevertheless, we use the source code to produce goal images by repeating the image generation several times. Some issues can still be observed to some extent, e.g., hallucination and loss of details. Moreover, both of these methods would hallucinate a robot or a human arm, originated from training data, which we believe may affect generalization.
Although the main contribution of this work is the goal pose prediction, we also demonstrate that a robot can indeed pick and place objects based on these goal poses. See our accompanying videos for demos. In the demo, Groundingdino [40], GraspNet [41] and MoveIt [42] are used for detection, grasping and motion planning, respectively.
V Conclusions
In this paper, we demonstrate the potential of achieving general-purpose object rearrangement with precise “something something” instructions. Specifically, given an instruction in natural language format, an LLM is used to identify the objects to be moved and to be the reference. Given RGB-D images, we utilize a pre-trained LLM-enhanced image segmentation model to obatin required objects and then point clouds. Based on these results, a diffusion-based 3D pose estimation model can follow precise low-level instructions to achieve physically-realistic position predictions. By establishing a GPT-assisted pipeline from a 3D perspective, a high-quality dataset for the placement task is generated. The results from both simulation and real-world experiments demonstrate the effectiveness of our approach. The model trained with simulation data can seamlessly transfer to real-world scenarios, achieving promising performance.